Evaluate the given definite integral. S (1-x)?dx dx = (Type an integer or a simplified fraction.)
Answer #1

44) 4. , 14 () 245 (44)) - 11 4 1 - 2 43 - แสน 4 - 2 ม. jo = (8 ) + (8) -2() (81) ๑) = 1 + (8)* 149 : 8) - 4s61 - (129) 5 4 1